I am presently thirty eight years old my son Chris is eleven years old. How many years ago was my age exactly four times his age?
Answer by rwm(914)

(38-x)=4*(11-x)
38-x=44-4x
3x=6
x=2
check
36=4*(9)
36=36
ok